A sample of carbon dioxide occupies 10.0 L under STP conditions.
At what pressure would this sample have a volume of 12.5 L and a temperature of 20.0 °C?
1
Expert's answer
2020-04-03T11:50:47-0400

n(CO2)=V(CO2)/Vm=10/22.4=0.45mol

Vm=22.4


PV=nRT

R=8.31; V=12.5l; T=20+273=293K


P*12.5=0.45*8.31*293

P*12.5=1095.7

P=87.7


Answer: 87.7